# references.bib

### v1

@@ 3 Methods and data
Citation counts for each publication are calculated using the full PMC Open Access Subset dataset (M publications above). Citations are based on identifiers, hence only references that include a valid ID are considered. Under these limits, we calculate total citation counts and use this as our main dependent variable. We also calculate citations given within a certain time window from publication (1, 2, and 3 years, also considering the month of publication). This is done in order to conduct a robustness check using citation counts over the same citation accrual time as the dependent variable (e.g., the three-year window for a publication published in June 2015 runs to June 2018 excluded).
@@ 50 Philosophy and Religious Studies 399 division_18
The reputation of authors before publication has also been linked to the citation success of a paper [sekara_chaperone_2018]. To control for this, we have to identify individual authors, a challenging task in itself [torvik_author_2009, lu_pubmed_2011, ferreira_brief_2012, liu_author_2014, zheng_entity_2015]. We focus on a publication-level aggregated indicator of author popularity: the mean H-index of a publication’s authors at the time of publication, calculated from the PMC Open Access Subset. In so doing, we minimize the impact of errors arising from disambiguating author names [strotmann_author_2012, kim_distortive_2016], which would have been higher if we had used measures based on individual observations such as the maximum H-index. We therefore use a simple disambiguation technique when compared to the current state of the art, and consider two author mentions to refer to the same individual if both full name and surname are found to be identical within all of the PMC Open Access Subset. We acknowledge the limitations of this method in possibly merging different authors with the same name and surname. We identify 8’481’129 seemingly distinct authors in this way.
@@ 4.1 Overview of the Open Science Indicators dataset
As mentioned previously, the OSI dataset we use for analysis comprises 121’999 articles. The majority are articles published in PLOS journals, with the largest proportion originating from PLOS ONE. The remaining articles have been taken from 1232 different journals published by a range of publishers. Rates of adoption for each Open Science practice can be calculated from the dataset to give an overall impression of the degree to which Open Science is practiced. Table 4 outlines the overall rates of adoption for the main Open Science practices in the dataset and shows that data (in a repository) and code sharing have a relatively low adoption rate across the dataset.
@@
The prevalence of different Open Science practices varies by field of research (following the divisions presented in 1). For example, both Division 3 (Health Sciences) and Division 8 (Engineering) have the lowest data sharing rate at 60%, whilst Division 16 (Language, Communication and Culture) has the highest data sharing rate at 82%. Similar degrees in variation are seen for the other indicators with data sharing in a repository ranging from 14% to 43%, code sharing from 7% to 36%, and preprint posting from 10% to 33%. Such wide variation in OSI adoption across divisions suggest that research fields face different challenges in adopting Open Science practices, and some practices may not be equally useful or relevant across fields. In Figure 3, we show trends for the main OSIs across all Divisions, as described in Table 1.
@@ 4.2 Modeling
The base model results we discuss are provided in Table LABEL:tab:model_base. It contains the basic author, publication, and journal-level variables we discussed above. It does not contain publication-level division classification. The most complete model we discuss is instead provided in Table LABEL:tab:model_full. Here, we use all the previous variables from the base model and add the publication-level division classification as dummy variables (division 1 to 18, see Table 1. Several more models were tested, primarily as robustness checks, and are discussed in the Appendix.
The base model is described in Equation 1, and the full model is described in Equation 2. Variable transformations are shown, numerical variables are given in Italics, and categorical variables are in regular text. Variables are grouped along lines. An illustration of the assumed causal dependency graph among variable groups is given in Figure 4. In the same figure, the variables for which we used a log scaling to limit the effects of outliers are flagged as such. These include the dependent variable (n_cit_tot), which is always used on a log scale.
@@ 5.2 Extension of previous research
While previous work [vandewalle_code_2012, cao_rise_2023, kang_papers_2023] has found as much as a threefold citation advantage for code sharing, we did not confirm this finding in our sample. Following [escamilla_rise_2022], it is possible that outside of fields like computer science authors are more likely to cite or link to shared code directly rather than citing the research paper with which it was associated.
@@ 6 Conclusion
Three of the authors (LC, ML, and IH) were at the time of publication employed by PLOS, the publisher of PLOS ONE.

### v2

@@ 1 Introduction
Despite their limitations [dougherty_citation_2022], citation counts and other bibliometric indicators are frequently used as quantitative measures of research impact and quality, and play a role in research assessment activities that support researchers’ career progression, and awarding of research funding [aksnes_citations_2019]. As such, actions that can increase the potential for citations of researchers’ articles may be seen as desirable and may motivate changes in research and publishing practices. As monitoring of Open Science increases, there is also growing recognition that those who are subject to monitoring by indicators should be involved in the development of indicators [himanen_scope_2024]. Biomedical researchers have, for example, developed community consensus on the relative importance, for monitoring, of 19 different Open Science practices [cobey_community_2023].
@@ 3 Methods and data
Citation counts for each publication are calculated using the full PMC Open Access Subset dataset (M publications above). Citations are based on identifiers, hence only references that include a valid ID are considered. Citations accumulated by the preprint of a publication are therefore not counted, as the preprint will have a different identifier than the published paper as included in OSI. Citations from a preprint to a published paper that is part of OSI are, however, counted. Under these constraints, we calculate total citation counts and use this as our main dependent variable. We also calculate citations given within a certain time window from publication (1, 2, and 3 years, also considering the month of publication). This is done in order to conduct a robustness check using citation counts over the same citation accrual time as the dependent variable (e.g., the three-year window for a publication published in June 2015 runs to June 2018 excluded).
@@ 50 Philosophy and Religious Studies 399 division_18
The reputation of authors before publication has also been linked to the citation success of a paper [sekara_chaperone_2018]. To control for this, we have to identify individual authors, a challenging task in itself [torvik_author_2009, lu_pubmed_2011, ferreira_brief_2012, liu_author_2014, zheng_entity_2015]. We focus on a publication-level aggregated indicator of author productivity and popularity: the mean H-index of a publication’s authors at the time of publication, calculated from the PMC Open Access Subset. While we acknowledge that the H-index has limitations and may, for instance, generate inconsistent rankings [waltman_inconsistency_2012], in our work we do not use the H-index to this end but instead use it as a proxy for the productivity and popularity of the authors of a publication. By using the mean H-index of the authors of a publication, we further minimize the impact of errors arising from disambiguating author names [strotmann_author_2012, kim_distortive_2016], which would have been higher if we had used measures based on individual observations such as the maximum H-index. We therefore use a simple disambiguation technique when compared to the current state of the art, and consider two author mentions to refer to the same individual if both full name and surname are found to be identical within all of the PMC Open Access Subset. We acknowledge the limitations of this method in possibly merging different authors with the same name and surname. We identify 8’481’129 seemingly distinct authors in this way.
@@ 4.1 Overview of the Open Science Indicators dataset
As mentioned previously, the OSI dataset we use for analysis comprises 121’999 articles. The majority of the articles are published in PLOS journals, with the largest proportion originating from PLOS ONE. The remaining articles have been taken from 1232 different journals published by a range of publishers. Rates of adoption for each Open Science practice can be calculated from the dataset to give an overall impression of the degree to which Open Science is practiced. Table 4 outlines the overall rates of adoption for the main Open Science practices in the dataset and shows that data (in a repository) and code sharing have a relatively low adoption rate across the dataset.
@@
The prevalence of different Open Science practices varies by field of research (following the divisions presented in 1). For example, both Division 3 (Health Sciences) and Division 8 (Engineering) have the lowest data sharing rate at 60%, whilst Division 16 (Language, Communication and Culture) has the highest data sharing rate at 82%. Similar degrees in variation are seen for the other indicators with data sharing in a repository ranging from 14% to 43%, code sharing from 7% to 36%, and preprint posting from 10% to 33%. Such wide variation in OSI adoption across divisions suggests that research fields face different challenges in adopting Open Science practices, and some practices may not be equally useful or relevant across fields. In Figure 3, we show trends for the main OSIs across all Divisions, as described in Table 1.
@@ 4.2 Modeling
The base model results we discuss are provided in Table LABEL:tab:model_base. It contains the basic author, publication, and journal-level variables we discussed above. It does not contain publication-level division classification. The most complete model we discuss is instead provided in Table LABEL:tab:model_full. Here, we use all the previous variables from the base model and add the publication-level division classification as dummy variables (division 1 to 18, see Table 1). Several more models were tested, primarily as robustness checks, and are discussed in the Appendix.
The base model is described in Equation 1, and the full model is described in Equation 2. Variable transformations are shown, numerical variables are given in Italics, and categorical variables are in regular text. Variables are grouped along lines. An illustration of the assumed causal dependency graph among variable groups is given in Figure 4. In the same figure, the variables for which we used log scaling to limit the effects of outliers are flagged as such. These include the dependent variable (n_cit_tot), which is always used on a log scale.
@@
We further zoom in on individual divisions by running the base model using data points from each division only and do not find significant effects for code sharing. The effect of sharing data in an online repository varies, but is larger in two divisions (division 2, Biological Sciences, and division 8, Engineering). The effect of preprints remains consistent across divisions, with varying degrees of significance and magnitude.
@@ 5.2 Extension of previous research
While previous work [vandewalle_code_2012, cao_rise_2023, kang_papers_2023] has found as much as a threefold citation advantage for code sharing, we did not confirm this finding in our sample. Following [escamilla_rise_2022], it is possible that outside of fields like computer science authors are more likely to cite or link to shared code directly rather than citing the research paper with which it was associated. Another possible reason could be that the quality of the description of the code, rather than its availability, impacts whether or not a research paper is cited, as was suggested in previous work on model papers [janssen_code_2020]. Without further research into code sharing and citation practices it is difficult to explain why there is a lack of significant findings for code.
@@ 6 Conclusion
Three of the authors (LC, ML, and IH) were at the time of publication employed by PLOS, the publisher of PLOS ONE. This does not alter our adherence to PLOS ONE policies on sharing data and materials.